Shift report 20020107 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Jan 07 2002 - 15:57:07 EST

  • Next message: Apache: "Shift report 20020107 16:00-24:00"

    ------------------------------------------------------------
    Date: Mon Jan  7 15:57:07 2002, Shift: 08:00-16:00
    Supervisor: E.J. Kim, with: Y.K. Lee
    Log book#20 pages 15 - 20
    
    Shift Summary:
    ------------------------------------------------------------
    
    1. Beam for data-run was lasted about 2hours.
       Polarization studies, deceleration, QLI etc.
       Now "Machine Setup & working on injectin"
    
    2. Data run 6331 : 248k trigger 3   Pedestal & DVM : run 6337, 6338, 6339
    
    3. Same problem with previous shift.
        --> D1D1, D1U2 - tripped (unable to reset)
        --> BBL41, BBL43, BBL44 - tripped (unable to reset)
        --> Need to access !
    
    ----------------------
       FV modified trigger (narrow gate ~ 6nsec overlap etc)
       FV & Kris changed "badtrip" to detect FS/MRS trigger
       counter trips.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Jan 07 2002 - 15:58:05 EST